OpenStudy (anonymous):

NEED HELP SOLVING 4X SQUARE= 12X-2

OpenStudy (anonymous):

IM SORRY -12

OpenStudy (anonymous):

do you mean \[4x ^{2}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

\[4x ^{2}+12x+2=0\]

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[\left( 2x+? \right)(2x+?)\]

OpenStudy (anonymous):

it should be -12

OpenStudy (anonymous):

i know you added it ti the other side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how to foil First inner outer last

OpenStudy (anonymous):

i do but his problem uses the quadratic formula to solve the equation

OpenStudy (anonymous):

\[\frac{ -b+or-\sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

\[ax ^{2}+bx+c\]

OpenStudy (anonymous):

a=4 b=12 c=2

OpenStudy (anonymous):

CORRECT

OpenStudy (anonymous):

can you plug in the numbers

OpenStudy (anonymous):

STAND BY

OpenStudy (anonymous):